Shift report 20030208 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 08 2003 - 00:07:34 EST

  • Next message: Apache: "Shift report 20030208 00:00-08:00"
    ------------------------------------------------------------
    Date: Sat Feb  8 00:07:34 2003, Shift: 16:00-24:00
    Supervisor: YKL, with: IA
    Log book#14 pages 4 - 4
    
    Shift Summary:
    ------------------------------------------------------------
    1. We had production beam for 2.5 hours. They lost the beam   
       because of a quenching link interlock.
    2. Data runs : <br>
       7829 MRS 40 deg. A500 100k; FS 8 deg. 1/15B 56k <br>
       7830 MRS 45 deg. A500 100k; FS 8 deg. 1/15B 81k <br>
       7831 MRS 90 deg. A350 2k; FS 8 deg. 1/15B 4k <br>
       7832 MRS 90 deg. A350 5k; FS 8 deg. 1/4B 4k <br>
    3. No problem with the detectors during our shift, but PS put a note regarding T3 so you should read it; or read his shift Elog.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Feb 08 2003 - 00:11:24 EST